# Instance Generation for Meta-Black-Box Optimization through Latent Space Reverse Engineering

### v1

@@ 3.3 Diverse-BBO
With the latent instance space and the GP searching strategy at hand, we now elaborate on how LSRE generates uniform-diversity function instances for Diverse-BBO. We first demarcate a grid area in the 2-dimensional latent space, i.e., $x\in[-B,B]$ and $y\in[-B,B]$, which could cover the target problem range $\mathcal{D}$ we constructed in Sec. 3.1. Then we uniformly sample $K\times K$ latent instance points (point matrix) from this grid area, where $K$ controls the sampling granularity. At last, for each sampled instance point, we call our GP search to reverse engineer a proximal symbolic tree and interpret it into the corresponding function formulation. In this paper, we set $B=1$ and $K=16$ to obtain a problem set of 256 diversified instances, which we denote as Diverse-BBO.
@@
To obtain Diverse-BBO is challenging considering that: 1) we have to repeatedly call the GP process for 256 times; 2) Within each GP process, evaluating the tree population is also a tedious and time-consuming task due to repeated objective computation. To accelerate LSRE, we further introduce Ray [^3] https://github.com/ray-project/ray to first distribute 256 GP processes on independent CPU cores and then trigger second-level parallel evaluation within each GP process. By doing this, we reduce the time cost of generating Diverse-BBO from days to hours.
We leave the hyper-parameters including their selection standards and the detailed design philosophy behind LSRE’s components at Appendix A.5 and Appendix A.6.
4 Experimental Analysis
[…]
  • Diverse-BBO: it includes 256 diversified synthetic problems generated by our LSRE, the problem dimensions for these instances are 2, 5, and 10.
  • CoCo-BBOB (Hansen et al. 2010a): it comprises 24 basis synthetic functions, which we used to generate 256 10-dimensional instances by transforming them via random rotation and shift.
  • MA-BBOB (Vermetten et al. 2025): it uses 24 basis fucntions in CoCo-BBOB as component functions and generates new instance by affine combination of them. We use MA-BBOB to generate 256 10-dimensional instances as a baseline.
  • GP-BBOB (Muñoz and Smith-Miles 2020): it is the first framework that uses GP for BBO instance generation. Our LSRE differs with this method in instance space analysis, symbol set construction and searching strategy. We use it to generate 256 instances as a baseline.
@@ 4.1 Diversity Comparison (RQ1)
Figure 4: Average Success Rate Distribution
We first show how we measure the diversity of a benchmark. Existing literature of BBO benchmarks barely provide explicit definition of diversity for the testing problem instances included. However, some existed papers indeed provide indirect measurements that reflect benchmark diversity through algorithm performance distribution (Hansen et al. 2010b). Following this idea, we propose a simple and intuitive average success rate distribution (ASRD) test to reflect the diversity of a benchmark through the problem difficulty distribution. In ASRD test, for a given benchmark that includes $N$ problem instances $f_{1}\sim f_{N}$, we first construct a BBO optimizer pool that comprises $M$ representative optimizers $\mathcal{A}_{1}\sim\mathcal{A}_{M}$. For any pair of problem instance $f_{i}$ and optimizer $\mathcal{A}_{j}$, we use $\mathcal{A}_{j}$ to optimize $f_{i}$ for 51 independent runs and record the portion of successful runs (runs that achieve optimal on $f_{i}$) as success rate $SR_{i,j}$. Then the ASRD for the given benchmark is summarized as the statistical histogram of all success rate results $\{\{SR_{i,j}\}_{i=1}^{N}\}_{j=1}^{M}$.
In this experiment, we conduct ASRD test for all four baselines, the optimizer pool we construct include several representative optimizers such as DE (Storn and Price 1997), PSO (Kennedy and Eberhart 1995), CMA-ES (Beyer and Schwefel 2002) etc.. For more detailed experiment protocol, we kindly request readers refer to Appendix B.1. We visualize the test results in Fig. 4, which show that: 1) Diverse-BBO v.s. CoCo-BBOB: derived from the latent space of CoCo-BBOB, our Diverse-BBO achieves preferred problem diversity through fine-grained generation of LSRE; 2) Diverse-BBO v.s. MA-BBOB: certain diversity improvement can be observed in MA-BBOB compared to CoCo-BBOB, however, the affine combination in MA-BBOB still fails to achieve controllable uniform diversity distribution; 3) Diverse-BBO v.s. GP-BBOB: sharing the same GP-based generation backbone, the novel designs we proposed in LSRE result in better generation quality, which we will discuss more in the subsequent sections.
@@ 4.2 Generalization Test (RQ2)
For the four baseline benchmarks, we use them for training MetaBBO approaches and then observe their generalization performance. This results in three experiment protocols we have to clarify here. First, for each baseline, we randomly split it into 75% training instances and 25% testing instances. We then collect all testing instances of the four benchmarks into one test set, termed as $\mathbb{D}_{synthetic}$. Second, we further prepare another three test sets $\mathbb{D}_{hpo}$, $\mathbb{D}_{uav}$ and $\mathbb{D}_{protein}$, which denote complex realistic BBO problems: hyper-parameter optimization (Arango et al. 2021), uav planning (Shehadeh and Kudela 2025) and protein docking (Hwang et al. 2010). Third, we select four representative MetaBBO approaches from existing literature: DEDDQN (Sharma et al. 2019), LDE (Sun et al. 2021), SYMBOL (Chen et al. 2024) and GLEET (Ma et al. 2024a). These baselines cover different methodology categories such as those for operator selection (DEDDQN), for algorithm configuration (LDE & GLEET) and for algorithm generation (SYMBOL). We use their default settings for training and testing according to the original papers. More details about these protocols are provided at Appendix B.2.
@@
We train the four MetaBBO approaches on the four baseline benchmarks’ train sets respectively, with the same random seed to ensure the comparison fairness. We then test these 16 (4 MetaBBOs, 4 train sets) trained models on $\mathbb{D}_{synthetic}$, with 51 independent runs on each problem instance. We adopt the average normalized optimization performance metrics (larger is better) widely used in MetaBBO approaches (Guo et al. 2025b; Ma et al. 2025b; Li et al. 2024) in the test. The testing results are presented in Table 1, where the ranks of baseline benchmarks are annotated alongside the performance metric, and the average ranks and performance across diverse MetaBBOs are also presented in the rightmost column. The results show that generally the diversity advantage of our Diverse-BBO against the other baselines facilitates the generalization performance of various MetaBBO approaches. It only slightly underperforms MA-BBOB with large overlapping. An interesting observation is that though MA-BBOB and GP-BBOB have been claimed useful for evaluating traditional BBO optimizer, they show an opposite effect in MetaBBO domain, which anticipates further investigation in the future.
@@
According to the insights obtained from recent MetaBBO-related researches, out-of-distribution generalization, especially MetaBBO’s zero-shot performance on unseen realistic problems should be thoroughly evaluated (Ma et al. 2025c). To this end, in this section we set a very high bar for the four baseline benchmarks. We test the four MetaBBO approaches trained previously on three realistic problem sets $\mathbb{D}_{hpo}$, $\mathbb{D}_{uav}$ and $\mathbb{D}_{protein}$. We illustrate in Fig. 5 the average performance gains of baseline benchmarks across the four MetaBBO approaches, which is computed as $\frac{1}{rank}$. The results further demonstrate the diversity advantage of our Diverse-BBO, indicating the effectiveness of LSRE in generating high-quality training instances for MetaBBO. To further interpret the observed superiority, we provide an intuitive visualization in Fig. 6 where the shaded red area denotes the coverage of all instances of the three realistic sets in the latent space $\mathcal{H}$ we constructed in Sec. 3.1. The blue points denotes the positions of the 256 instances of each baseline benchmark in $\mathcal{H}$. We believe results there could clearly explain the relationship between benchmark diversity and corresponding generalization potential. Through our LSRE, we achieve controllable, uniform and fine-grained instance generation within the instance space, which hence benefits training of MetaBBO.
@@ 4.3 Ablation Study (RQ3)
Figure 7: LSRE optimization curves when setting vary.
We provide following ablation baselines to validate our proposed novel designs in LSRE:
  • LSRE-PCA: where we use PCA for the latent instance space analysis, selecting the most important two features for dimension reduction.
  • LSRE-Simple: where we use simpler symbol set design of existing symbolic regression research (Tian et al. 2020), without vectorized operations.
  • LSRE-NO_LS: where the cross-dimensional local search strategy is ablated from the GP search process.
Besides, we additionally include a background baseline GP-BBOB (Muñoz and Smith-Miles 2020), which can be regarded as a baseline that ablates all novel designs in LSRE. Specifically, GP-BBOB share the same GP backbone as our LSRE, while it uses the simpler symbol set, the PCA-based instance space analysis and naive GP search without local search. We present in Fig. 7 the optimization curves of these ablation baselines and our LSRE when searching for $\tau^{*}$, which are averaged across all 256 target instance points. The results not only demonstrate significant improvement LSRE achieved against GP-BBOB, but also validate that each novel design in LSRE contributes to the final performance. Specifically, while the augmented symbol set and refined autoencoder-based instance analysis provide comparably equal contribution, we observe that the cross-dimensional local search plays a more important role in LSRE. Compared to existing works that fix the problem dimension during their searching process (Muñoz and Smith-Miles 2020; Long et al. 2023; Vermetten et al. 2025), our finding reveals that the ELA feature space might not be smooth enough if we do not co-optimize the problem dimension. There are also several insightful and interesting finds that can not be presented here due to the space limitation, e.g., the visualized comparison between PCA and autoencoder in instance space analysis, the complexity of the function instance generated by LSRE etc.. We leave them at Appendix B.3 for curious readers.

### v2

@@ 3.3 Diverse-BBO
With the latent instance space and the GP searching strategy at hand, we now elaborate on how LSRE generates uniform-diversity function instances for Diverse-BBO. We first demarcate a grid area in the 2-dimensional latent space, i.e., $x\in[-B,B]$ and $y\in[-B,B]$, which could cover the target problem range $\mathcal{D}$ we constructed in Sec. 3.1. Then we uniformly sample $K\times K$ latent instance points (point matrix) from this grid area, where $K$ controls the sampling granularity. At last, for each sampled instance point, we call our GP search to reverse engineer -corresponding function formulation. In this paper, we set $B=1$ and $K=16$ to obtain a problem set of 256 diversified instances, which we denote as Diverse-BBO. Refer to “‘DiverseBBO.pdf” in our codebase for their details.
@@
To obtain Diverse-BBO is challenging considering that: 1) we have to repeatedly call the GP process for 256 times; 2) Within each GP process, evaluating the tree population is also a tedious and time-consuming task due to repeated objective computation. The overall complexity is hence $\mathcal{O}[M*G*(N^{2}+N*L^{2})]$, where M is the number of functions we need to generate (256), G is the GP generations (50), N and L are population size and tree node size respectively. To accelerate LSRE, we further introduce Ray [^4] https://github.com/ray-project/ray to first distribute 256 GP processes on independent CPU cores and then trigger second-level parallel evaluation within each GP process. By doing this, we significantly reduce the overall complexity to $\mathcal{O}[G*L^{2})]$. We leave the hyper-parameters including their selection standards and the detailed design philosophy behind LSRE’s components at Appendix A.5 and Appendix A.6.
Figure 4: Average Success Rate Distribution
@@ 4.1 Diversity Comparison (RQ1)
We first show how we measure the diversity of a benchmark. Existing literature of BBO benchmarks barely provide explicit definition of diversity for the testing problem instances included. However, some existed papers indeed provide indirect measurements that reflect benchmark diversity through algorithm performance distribution (diversity-1). Following this idea, we propose a simple and intuitive average success rate distribution (ASRD) test to reflect the diversity of a benchmark through the problem difficulty distribution. In ASRD test, for a given benchmark that includes $N$ problem instances $f_{1}\sim f_{N}$, we first construct a BBO optimizer pool that comprises $M$ representative optimizers $\mathcal{A}_{1}\sim\mathcal{A}_{M}$. For any pair of problem instance $f_{i}$ and optimizer $\mathcal{A}_{j}$, we use $\mathcal{A}_{j}$ to optimize $f_{i}$ for 51 independent runs and record the portion of successful runs (runs that achieve optimal on $f_{i}$) as success rate $SR_{i,j}$. Then the ASRD for the given benchmark is summarized as the statistical histogram of all success rate results $\{\{SR_{i,j}\}_{i=1}^{N}\}_{j=1}^{M}$.
In this experiment, we conduct ASRD test for all four baselines, the optimizer pool we construct include several representative optimizers such as DE, PSO, CMA-ES etc.. For more detailed experiment protocol, we kindly request readers refer to Appendix B.1. We visualize the test results in Fig. 4, which show that: 1) Diverse-BBO v.s. CoCo-BBOB: derived from the latent space of CoCo-BBOB, our Diverse-BBO achieves preferred problem diversity through fine-grained generation of LSRE; 2) Diverse-BBO v.s. MA-BBOB: certain diversity improvement can be observed in MA-BBOB compared to CoCo-BBOB, however, the affine combination in MA-BBOB still fails to achieve controllable uniform diversity distribution; 3) Diverse-BBO v.s. GP-BBOB: sharing the same GP-based generation backbone, the novel designs we proposed in LSRE result in better generation quality, which we will discuss more in the subsequent sections.
@@ 4.2 Generalization Test (RQ2)
For the four baseline benchmarks, we use them for training MetaBBO approaches and then observe their generalization performance. This results in three experiment protocols we have to clarify here. First, for each baseline, we randomly split it into 75% training instances and 25% testing instances. We then collect all testing instances of the four benchmarks into one test set, termed as $\mathbb{D}_{synthetic}$. Second, we further prepare another three test sets $\mathbb{D}_{hpo}$, $\mathbb{D}_{uav}$ and $\mathbb{D}_{protein}$, which denote complex realistic BBO problems: hyper-parameter optimization (hpob), uav planning (uav) and protein docking (proteindocking). Third, we select four representative MetaBBO approaches from existing literature: DEDDQN (DEDDQN), LDE (LDE), SYMBOL (Symbol) and GLEET (GLEET). These baselines cover different methodology categories such as those for operator selection (DEDDQN), for algorithm configuration (LDE & GLEET) and for algorithm generation (SYMBOL). We use their default settings for training and testing according to the original papers. More details about these protocols are provided at Appendix B.2.
@@
We train the four MetaBBO approaches on the four baseline benchmarks’ train sets respectively, with the same random seed to ensure the comparison fairness. We then test these 16 (4 MetaBBOs, 4 train sets) trained models on $\mathbb{D}_{synthetic}$, with 51 independent runs on each problem instance. We adopt the average normalized optimization performance metrics (larger is better) widely used in MetaBBO approaches (ConfigX; neurela; GLHF) in the test. The testing results are presented in Table 1, where the ranks of baseline benchmarks are annotated alongside the performance metric, and the average ranks and performance across diverse MetaBBOs are also presented in the rightmost column. The results show that generally the diversity advantage of our Diverse-BBO against the other baselines facilitates the generalization performance of various MetaBBO approaches. It only slightly underperforms MA-BBOB with large overlapping. An interesting observation is that though MA-BBOB and GP-BBOB have been claimed useful for evaluating traditional BBO optimizer, they show an opposite effect in MetaBBO domain, which anticipates further investigation in the future.
@@
According to the insights obtained from recent MetaBBO-related researches, out-of-distribution generalization, especially MetaBBO’s zero-shot performance on unseen realistic problems should be thoroughly evaluated (metabox-v2). To this end, in this section we set a very high bar for the four baseline benchmarks. We test the four MetaBBO approaches trained previously on three realistic problem sets $\mathbb{D}_{hpo}$, $\mathbb{D}_{uav}$ and $\mathbb{D}_{protein}$. We illustrate in Fig. 5 the average performance gains of baseline benchmarks across the four MetaBBO approaches, which is computed as $\frac{1}{rank}$. The results further demonstrate the diversity advantage of our Diverse-BBO, indicating the effectiveness of LSRE in generating high-quality training instances for MetaBBO. To further interpret the observed superiority, we provide an intuitive visualization in Fig. 6 where the shaded red area denotes the coverage of all instances of the three realistic sets in the latent space $\mathcal{H}$ we constructed in Sec. 3.1. The blue points denotes the positions of the 256 instances of each baseline benchmark in $\mathcal{H}$. We believe results there could clearly explain the relationship between benchmark diversity and corresponding generalization potential. Through our LSRE, we achieve controllable, uniform and fine-grained instance generation within the instance space, which hence benefits training of MetaBBO.
@@
Figure 7: LSRE optimization curves when setting vary.
@@ 4.3 Ablation Study (RQ3)
We provide following ablation baselines to validate our proposed novel designs in LSRE: 1) LSRE-PCA: where we use PCA for the latent instance space analysis, selecting the most important two features for dimension reduction; 2) LSRE-Simple: where we use simpler symbol set design of existing symbolic regression research (GP-BBOB-3), without vectorized operations; 3) LSRE-NO_LS: where the cross-dimensional local search strategy is ablated from the GP search process. Besides, we additionally include a background baseline GP-BBOB (GP-BBOB-1), which can be regarded as a baseline that ablates all novel designs in LSRE. Specifically, GP-BBOB share the same GP backbone as our LSRE, while it uses the simpler symbol set, PCA-based instance space analysis and naive GP search.
We present in Fig. 7 the optimization curves of these ablation baselines and our LSRE when searching for $\tau^{*}$, which are averaged across all 256 target instance points. The results not only demonstrate significant improvement LSRE achieved against GP-BBOB, but also validate that each novel design in LSRE contributes to the final performance. Specifically, while the augmented symbol set and refined autoencoder-based instance analysis provide comparably equal contribution, we observe that the cross-dimensional local search plays a more important role in LSRE. Compared to existing works that fix the problem dimension during their searching process (GP-BBOB-1; GP-BBOB-2; ma-bbob), our finding reveals that the ELA feature space might not be smooth enough if we do not co-optimize the problem dimension. There are also several insightful and interesting findings such as the visualized comparison between PCA and autoencoder in instance space analysis, the complexity of the function instance generated by LSRE etc.. We leave them at Appendix B.3 for curious readers.